OpenStudy (anonymous):

Someone please help? Medal and Fan (: If A = (1, 4) B = (3, 1) and C = (6, 3), classify the the following triangle as either equilateral or right. Type the word equilateral or right in the blank.

OpenStudy (abb0t):

try calculating the length of the 3 sides if they are equal, it would be equilateral if not, you have to check if they apply Pythagoras theorem  Pythagorean theorem: \(\sf \color{red}{a^2+b^2=c^2}\)

OpenStudy (abb0t):

OR you can plot it! 

OpenStudy (anonymous):

I plotted it. I'm just confused on where the numbers fall into place @abb0t

OpenStudy (abb0t):

find the slope of the two lines that form the right angle and see if the slopes are negative reciprocals

OpenStudy (anonymous):

plot it in a graph...and calculate d all sides in the graph..

OpenStudy (anonymous):

I count the boxes?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Really? I literally just count the boxes in the shape? that would be my answer?

OpenStudy (anonymous):

i don't think counting boxes works in this case i think you need the distance formula

OpenStudy (anonymous):

all 3 sides are equal???

OpenStudy (anonymous):

distance between (1, 4) and (3, 1) is \(\sqrt{2^2+3^2}=\sqrt{13}\)

OpenStudy (anonymous):

Its a right triangle. am I correct?

OpenStudy (anonymous):

distance between (3, 1) and (6, 3) is \(\sqrt{3^2+2^2}=\sqrt{13}\) also

OpenStudy (anonymous):

distance between (1,4) and (3,6) is \(\sqrt{2^2+2^2}=\sqrt{8}\)

OpenStudy (anonymous):

so no, it is not equilateral

OpenStudy (anonymous):

and no, it is not a right triangle either

OpenStudy (anonymous):

if all 3 sides are equal den that is equilateral

OpenStudy (anonymous):

I don't understand... it could only be those two answers.. it makes more sense for it to be right triangle because of how it looks on the graph..

OpenStudy (anonymous):

are you saying that any triangle is either a right triangle or an equilateral triangle? that is certainly not the case this one is neither

OpenStudy (anonymous):

it is isosceles though, two sides are equal

it doesn't give me that option ):